# Cycle Matroids of Graphings: From Convergence to Duality

### v2

@@ 3.2 Non-continuity for local convergence
Finally, we prove Theorem 1.2, showing that local convergence does not guarantee quotient-convergence of the matroids.
Proof of Theorem 1.2..
Let $(G_{1},G_{2},\dots)$ be a sequence of $3$-regular, $c$-expander graphs with $|V(G_{n})|=2n$, such that their girth $g_{n}$ tends to infinity. Let $\mathcal{G}$ be a subsequential local-global limit of $G_{n}$. Note that $\mathcal{G}$ is a 3-regular treeing. Let $H_{n}=G_{n}\cup G_{n}$, that is, $H_{n}$ is the disjoint union of two copies of $G_{n}$. Clearly, the sequence $H_{n}$ locally converges to $\mathcal{G}$ as well. Write $\rho_{G_{n}}$ for $r_{G_{n}}/|V(G_{n})|$, and similarly for $H_{n}$.
The partition of $H_{n}$ into two copies of $G_{n}$ defines a factor of the cycle matroid:
$a_{n}\in Q_{2}(H_{n})\subseteq\mathbb{R}^{2^{2}},\ a_{n}(\emptyset)=0,\ a_{n}(\{1\})=a_{n}(\{2\})=\frac{|V(G_{n})|-1}{2|V(G_{n})|},\ a_{n}(\{1,2\})=\frac{2|V(G_{n})|-2}{2|V(G_{n})|}.$
@@
[…]
By the definition of expander sequences, for every $S\subseteq V(G_{n})$ with $|S|\leq n$, the number of edges connecting $S$ to $V(G_{n})\setminus S$ is at least $c\cdot|S|$. Let $E(G_{n})=E_{1}\cup E_{2}$ be any 2-partition of the edges. For $i\in\{0,1,2,3\}$, we denote by $V_{i}$ the set of vertices of $G_{n}$ with degree $i$ in $E_{1}$ and let $m_{i}=|V_{i}|$. Suppose that this partition results in a quotient closer to $a$ than $\varepsilon$. Then, $|\rho_{G_{n}}(E_{i})-\frac{1}{2}|<\varepsilon$ for $i=1,2$. Let us note that $r_{G_{n}}(E_{1})\leq 2n-m_{0}$, and the previous two inequalities imply $m_{0}/(2n)\leq(1/2+\varepsilon)$. The same bound holds for $m_{3}$.
We get a contradiction by showing that this is not possible if $g_{n}\geq 32/c$ and $\varepsilon<c/32$. We bound the number of $E_{1}$-components as follows. The number of singletons is clearly $m_{0}$. If an $E_{1}$-component is not a tree, then it has at least $g_{n}$ vertices, and hence the number of such components is at most $2n/g_{n}$. Finally, if a component is a non-singleton tree, then it has at least $2$ vertices off degree $1$, so the number of such components is bounded by $m_{1}/2$. We get that
$\rho_{G_{n}}(E_{1})\geq 1-\frac{1}{2n}\left(m_{0}+\frac{m_{1}}{2}+\frac{2n}{g_{n}}\right).$
Writing up the analogous inequality for $E_{2}$ and adding up the two inequalities, we get
$\rho_{G_{n}}(E_{1})+\rho_{G_{n}}(E_{2})\geq 2-\frac{1}{2n}\left(m_{0}+m_{3}+\frac{m_{1}+m_{2}}{2}\right)-\frac{2}{g_{n}}=1+\frac{m_{1}+m_{2}}{4n}-\frac{2}{g_{n}}.$ (3)
Without loss of generality, we may assume that $m_{0}\leq m_{3}$, implying $m_{0}\leq n$. The neighbours of $V_{0}$ in $G$ are contained in $V_{1}\cup V_{2}$, and hence
$m_{1}+m_{2}\geq cm_{0}.$
We also have
$m_{1}+m_{2}=2n-m_{0}-m_{3}\geq 2n-m_{0}-\left(\frac{1}{2}+\varepsilon\right)2n=\left(\frac{1}{2}-\varepsilon\right)2n-m_{0}.$
These two inequalities imply
$m_{1}+m_{2}\geq\frac{c}{c+1}\left(\frac{1}{2}-\varepsilon\right)2n.$
Thus, by (3) and the assumptions $g_{n}\geq 32/c$ and $\varepsilon<c/32$, we get
$\rho_{G_{n}}(E_{1})+\rho_{G_{n}}(E_{2})\geq 1+\frac{c}{2c+2}\left(\frac{1}{2}-\varepsilon\right)-\frac{2}{g_{n}}>1+2\varepsilon,$
contradicting $|\rho_{G_{n}}(E_{i})-\frac{1}{2}|<\varepsilon$ for $i=1,2$. ∎
@@ 4 Exposed minorizing measures and hyperfinite spanning forests
In this section, we focus on characterizing the exposed points of the base-polytope of the cycle matroid of a graphing. Recall that for a finite matroid $M=(E,r)$, the convex hull of its bases is described as
$C_{M}=\left\{x\in\mathbb{R}^{E}~{}\Big{|}~{}x\geq 0,\ \sum_{e\in A}x(a)\leq r(A)\text{ for }A\subseteq E,\ \sum_{e\in E}x(e)=r(E)\right\}.$
Let $\mathcal{G}=(X,E,\mu)$ be a graphing, let $\tilde{\mu}$ denote its edge measure, and $\rho$ the corresponding rank function. The analytic counterpart of $C_{M}$ is the convex set of basic minorizing finitely additive measures
$C_{\mathcal{G}}=bmm^{+}(\rho)=\big{\{}\alpha\textrm{ finitely additive measureon $E$}~{}\big{|}~{}\alpha\leq\rho,\ \alpha(E)=\rho(E)\big{\}}.$
In this section we write $C=C_{\mathcal{G}}$ for brevity. If $\mathcal{G}$ has no infinite components, then $\rho(E)=1$. Furthermore, as $\alpha\leq\rho\leq\tilde{\mu}$, in fact all $\alpha\in C$ are $\sigma$-additive measures and they admit Radon-Nikodym derivatives with respect to $\tilde{\mu}$, denoted by
$f_{\alpha}(x)=\frac{\partial\alpha}{\partial\tilde{\mu}}(x).$
Note that $f_{\alpha}$ takes values in $[0,1]$. If $f_{\alpha}$ is 00-$1$ valued $\tilde{\mu}$-almost everywhere, then $\alpha$ is clearly extreme in $C$. In this case there is a corresponding measurable set of edges $E_{\alpha}=\{e\in E\mid f_{\alpha}(e)=1\}\subseteq E$. The first observation we make is that such $\mathcal{T}=(X,E_{\alpha},\mu)$ is an essential hyperfinite spanning forest.
@@
If $f_{\alpha}$ is 0-1 valued, then $\mathcal{T}=(X,E_{\alpha},\mu)$ is an essential hyperfinite spanning forest.
@@
By $\alpha\leq\rho$, $E_{\alpha}$ contains no cycles, so $\mathcal{T}$ is a forest. As $\alpha\in\text{bmm}^{+}(\rho_{\mathcal{G}})$, we have $\rho_{\mathcal{G}}(E_{\alpha})=\rho_{\mathcal{G}}(E)$, and therefore $\mathcal{T}$ is essential. Also, we have $\textrm{cost}(\mathcal{T})=\tilde{\mu}(E_{\alpha})=\alpha(E)=\rho_{\mathcal{T}}(E_{\alpha})$, so $\mathcal{T}$ is hyperfinite by Proposition 2.2. ∎
For an essential hyperfinite spanning forest $\mathcal{T}=(X,F,\mu)$, let $\alpha_{F}$ denote the measure defined by $\alpha_{F}(A)=\tilde{\mu}(A\cap F)$. We have the following.
Lemma 4.2.
For any essential hyperfinite spanning forest $\mathcal{T}=(X,F,\mu)$ of a graphing $\mathcal{G}$, $\alpha_{F}$ is in $C$.
Proof.
Let $A\subseteq E$ be a Borel subset of the edges of $\mathcal{G}$. Then $\mathcal{T}_{A}=(X,A\cap F,\mu)$ is a hyperfinite forest and $\tilde{\mu}(A\cap F)=\textrm{cost}(\mathcal{T}_{A})$. By Proposition 2.2, we get $\textrm{cost}(\mathcal{T}_{A})=\rho(A\cap F)$. By monotonicity of $\rho$, this implies $\alpha_{F}(A)\leq\rho(A)$. ∎
Remark 4.3.
Observe that in the above argument, the equality $\alpha_{F}(A)=\rho(A)$ holds if and only if $\rho(A\cap F)=\rho(A)$. This happens exactly when for $\mu$-almost all $x\in X$ the $A$- and $(A\cap F)$-components of $x$ have the same size. In case this size is infinity, the $(A\cap F)$-component does not need to be the whole $A$-component. In other words, there is equality exactly when $\mathcal{T}_{A}=(X,A\cap F,\mu)$ is essential in $\mathcal{G}_{A}=(X,A,\mu)$.
4.1 Exposed points
We say that $f_{\alpha}$ is exposed by $g\in L^{1}(E,\tilde{\mu})$ if the corresponding continuous linear functional $\varphi_{g}$ of $L^{\infty}(E,\tilde{\mu})$ attains its strict maximum on $C$ at $f_{\alpha}$. More formally, if for any $\beta\in C$, $\beta\neq\alpha$ we have $\varphi_{g}(f_{\alpha})>\varphi_{g}(f_{\beta})$. Clearly, exposed points are extreme. We prove that all exposed points of $C$ are of the form $\alpha_{F}$, or in other words, all the exposed $f_{\alpha}$ are 00-$1$ valued.
Theorem 4.4.
Let $f_{\alpha}\in C$ be exposed by a function $g\in L^{1}(E,\tilde{\mu})$. Then, $f_{\alpha}$ is 00-$1$ valued $\tilde{\mu}$-almost everywhere.
Remark 4.5.
Usually, for a convex set $C$ in a topological vector space, a point $x\in C$ is called exposed if there exists a continuous linear functional that attains its strict maximum at $x$. In our case, $C\subseteq L^{p}(E,\tilde{\mu})$ for all $p\in[1,\infty]$. For $f_{\alpha}$ to be exposed in $L^{p}(E,\tilde{\mu})$ where $p\neq\infty$, the witness $g$ has to be in $L^{q}(E,\tilde{\mu})$, where $q$ is the dual pair of $p$, that is, $11/p+1/q=$ 1. This means that it is easier to be exposed for higher values of $p$ since there are more functions in $L^{q}$ for lower values of $q$. It is even easier to be exposed in $L^{\infty}(E,\tilde{\mu})$, even with the restriction that the linear functional corresponds to an $L^{1}$ function, and this is enough for our argument. What is not enough is to be exposed in $L^{\infty}(E,\tilde{\mu})$ in general, because the linear functional might correspond to a finitely additive measure, and not to a measure whose Radon-Nikodym derivative would be the witness $g\in L^{1}(E,\tilde{\mu})$.
Proof of Theorem 4.4..
We treat $g$ as a weight function on the edges, and our goal is to find a maximal essential hyperfinite spanning forest $F$ with respect to this weight function. We write $\alpha_{F}$ for the corresponding measure in $C$, i.e. $\alpha_{F}(A)=\tilde{\mu}(A\cap F)$ for all measurable $A\subseteq E$. The algorithm will result in an $F$ satisfying $\alpha_{F}(\{g\geq t\})=\rho(\{g\geq t\})$. In a layer-cake integration this implies, for all $\beta\in C$,
@@ 4.1 Exposed points
We run a version of Kruskal’s greedy algorithm [14] on $\mathcal{G}$. To handle ties, we fix an arbitrary Borel ordering of $E$ in advance, and whenever there is a tie in the $g$ values among finitely many edges considered, we choose the maximal one according to the Borel ordering. The algorithm constructs Borel edge sets $F_{1}\subseteq F_{2}\subseteq\ldots$, with $F=\bigcup F_{n}$ being the final output. In the first step, every vertex simultaneously chooses the incident edge with the highest $g$-value to be included in $F_{1}$. During the $n^{\mathrm{th}}$ step, every finite component of $F_{n-1}$ adds the edge with the highest $g$-value from the set of edges leaving it, and thus we obtain $F_{n}$. In case there is no edge leaving a finite $F_{n-1}$ component we do not add anything, but observe that in this component $F$ is already a spanning tree. Also note that infinite components do not add any edges, but it may happen that a finite component adds an edge that connects it to an infinite one.
@@
It is not difficult to see that the sets $F_{n}$ are Borel, cycles cannot be created, and $F$ is essential. We claim that $\tilde{\mu}(F)=\rho_{\mathcal{G}}(E)$, so $F$ is hyperfinite by Proposition 2.2. Indeed, set up a payment function as follows. In the beginning, every vertex receives $1$ token. In a general step, every finite connected component gives $1$ token to the edge it chooses. After the addition of the edges, for every finite connected component there is exactly one edge who received $2$ tokens. This edge gives $1$ token back to the component, thus ensuring that every finite component has a token before the next step. Should a finite $F$-component become spanning in a finite $\mathcal{G}$-component, and therefore has no use for its token, this token is redistributed equally among the vertices of this component. In the end, vertices in finite $\mathcal{G}$-components paid $1-\frac{1}{|\mathcal{G}_{x}|}$ token, and vertices in infinite $\mathcal{G}$-components paid 1 token. On the other hand every chosen edge received 1 token, hence $\rho_{\mathcal{G}}(E)=\tilde{\mu}(F)$ by the Mass Transport Principle. As $F$ is hyperfinite, by Lemma 4.2, we have $\alpha_{F}\in C$.
Let $A_{t}=\{g\geq t\}$ and $F_{\geq t}=F\cap A_{t}$. We further observe that $F_{t}$ is an essential spanning forest of $A_{t}$. That is, for $\mu$-almost every $x\in X$,
@@
By Remark 4.3, (1) and (2) imply $\alpha_{F}(A_{t})=\rho(A_{t})$, completing the proof of the theorem. ∎
@@
If $f_{\alpha}$ is 00-$1$ valued, then it is exposed – the witness for this is simply choosing $g=f_{\alpha}$. In fact, in this case $f_{\alpha}$ is strongly exposed, i.e. for any sequence $\beta_{n}\in C$ we have $\varphi_{g}(f_{\beta_{n}})\to\varphi_{g}(f_{\alpha})\Rightarrow||f_{\beta_{n}}-f_{\alpha}||_{p}\to 0$ for all $p\in[1,\infty)$. So Theorem 4.4 characterizes exposed points of $C$.
@@
By the Klee–Straszewicz theorem [13, Theorem 2.1], exposed points are dense among extreme points of $C$ in the weak topology of $L^{\infty}(E,\tilde{\mu})$. Moreover, the exposed points determine $C$ as $C=\overline{\textrm{conv}(\exp(C))}^{w^{*}}$.

### v3

@@ 3.2 Non-continuity for local convergence
It was proved in [23, Theorem 4.7] that the rank of the whole edge set of a graph is locally estimable, i.e. continuous with respect to local convergence [21, Theorem 22.3]. We show by an example that local convergence does not guarantee quotient-convergence of the matroids.
Example 3.4.
Let $(G_{1},G_{2},\dots)$ be a sequence of $3$-regular, $c$-expander graphs with $|V(G_{n})|=2n$ and $10<c<$ 1, such that their girth $g_{n}$ tends to infinity. Let $H_{n}=G_{n}\cup G_{n}$, that is, $H_{n}$ is the disjoint union of two copies of $G_{n}$. Clearly, the interlaced sequence $G_{1},H_{1},G_{2},H_{2},\dots$ is locally convergent (it converges to any 3-regular treeing). Let $\rho_{G_{n}}=r_{G_{n}}/|V(G_{n})|$, and similarly for $H_{n}$. We claim that $Q_{2}(\rho_{G_{n}})$ is far from $Q_{2}(\rho_{H_{m}})$ if $n$ and $m$ are large, so the sequence $(\rho_{G_{1}},\rho_{H_{1}},\rho_{G_{2}},\rho_{H_{2}},\dots)$ is not quotient-convergent.
The partition of $H_{n}$ into two copies of $G_{n}$ defines a quotient of the cycle matroid:
$a_{n}\in Q_{2}(H_{n})\subseteq\mathbb{R}^{2^{2}},\ a_{n}(\emptyset)=0,\ a_{n}(\{1\})=a_{n}(\{2\})=\frac{2n-1}{4n},\ a_{n}(\{1,2\})=\frac{4n-2}{4n}.$
@@
We show that if $n$ is sufficiently large, then $G_{n}$ cannot have $2$-quotients that are $\varepsilon$-approximations of $a$, where $\varepsilon=c/32$. This will imply that $d^{\text{\rm Haus}}(Q_{2}(\rho_{G_{n}}),Q_{2}(\rho_{H_{m}}))>\varepsilon$. We choose $n$ large enough so that $g_{n}>32/c$.
By the definition of expander sequences, for every $S\subseteq V(G_{n})$ with $|S|\leq n$, the number of edges connecting $S$ to $V(G_{n})\setminus S$ is at least $c\cdot|S|$. Let $E(G_{n})=E_{1}\cup E_{2}$ be any 2-partition of the edges. For $i\in\{0,1,2,3\}$, we denote by $V_{i}$ the set of vertices of $G_{n}$ with degree $i$ in $E_{1}$ and let $m_{i}=|V_{i}|$. Suppose that this partition results in a quotient closer to $a$ than $\varepsilon$. Then, $|\rho_{G_{n}}(E_{i})-\frac{1}{2}|<\varepsilon$ for $i=1,2$, and so $|r_{G_{n}}(E_{i})-n|<2\varepsilon n$. Let us note that $r_{G_{n}}(E_{1})\leq 2n-m_{0}$, and the previous two inequalities imply $m_{0}\leq(1+2\varepsilon)n$. The same bound holds for $m_{3}$.
We bound the number of $E_{1}$-components as follows. The number of singletons is clearly $m_{0}$. If an $E_{1}$-component is not a tree, then it has at least $g_{n}$ vertices, and hence the number of such components is at most $2n/g_{n}$. Finally, if a component is a non-singleton tree, then it has at least $2$ vertices of degree $1$, so the number of such components is bounded by $m_{1}/2$. Hence
$m_{0}+\frac{m_{1}}{2}\geq 2n-r_{G_{n}}(E_{1})-\frac{2n}{g_{n}}\geq\Big{(}1-2\varepsilon-\frac{2}{g_{n}}\Big{)}n\geq\Big{(}1-\frac{c}{8}\Big{)}n.$
Adding the analogous inequality for $E_{2}$, we get
$2m_{0}+m_{1}+m_{2}+2m_{3}\geq\Big{(}2-\frac{c}{4}\Big{)}n,$
and using that $m_{0}+m_{1}+m_{2}+m_{3}=2n$, we obtain $m_{1}+m_{2}\leq cn/4$. Without loss of generality, we may assume that $m_{0}\leq m_{3}$, implying that $m_{0}\leq|V(G_{n})|/2$. The neighbours of $V_{0}$ in $G$ are contained in $V_{1}\cup V_{2}$, and hence by the expansion of $G_{n}$, we have $m_{1}+m_{2}\geq cm_{0}$. So,
$m_{0}+m_{1}+m_{2}+m_{3}\leq\frac{c+1}{c}(m_{1}+m_{2})+m_{3}\leq\frac{c+1}{c}\,\frac{c}{4}\,n+\Big{(}1+\frac{c}{16}\Big{)}n=\Big{(}1+\frac{5c}{16}\Big{)}n<2n.$
This is a contradiction.
@@ 4 Exposed minorizing measures and hyperfinite spanning forests
In this section, we focus on the set of minorizing finitely additive measures of the rank function of a graphing. In Section 4.1, we establish a connection between the Radon-Nikodym derivatives of such measures and hyperfinite essential spanning forests. We characterize the exposed points of minorizing finitely additive measures in Section 4.2, and then introduce the notion of weakly exposed points in Section 4.3. Based on these observations, we prove Theorems 1.2 and 1.3 in Section 4.4. Finally, in Section 4.5, we give a conceptually new proof of the submodularity of the rank function $\rho_{\mathbf{G}}$ of a graphing $\mathbf{G}$ that might be of independent combinatorial interest.
4.1 Measures, functions and subsets
Let $\mathbf{G}=(J,\mu,E)$ be a graphing and $\rho_{\mathbf{G}}$ be the rank function of its cycle matroid. If $\mathbf{G}$ has no finite components, then $\rho_{\mathbf{G}}(E)=1$. Since any $\alpha\in\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ is a $\sigma$-additive measure, it admits a Radon-Nikodym derivative with respect to $\eta_{\mathbf{G}}$, denoted by
$f_{\alpha}(x)=\frac{\partial\alpha}{\partial\eta_{\mathbf{G}}}(x).$
For a function $f\in L^{1}(\eta_{\mathbf{G}})$, let $\eta_{f}$ denote the measure
$\eta_{f}(A)=\int_{A}f\mathop{}\!\mathrm{d}\eta_{\mathbf{G}}.$
By the above, we can identify the convex sets of measures $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ and $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ with convex sets of measurable functions. By a slight abuse of notation, we say that a function $f\in L^{1}(\eta_{\mathbf{G}})$ is in $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ or $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ if the corresponding measure $\eta_{f}$ is in $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ or $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$, respectively. Note that $f_{\alpha}$ takes values in $[0,1]$.
If $f_{\alpha}$ is 00-$1$ valued $\eta_{\mathbf{G}}$-almost everywhere, then $\alpha$ is clearly extreme in both $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ and $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$. In such a case, let us consider the corresponding measurable set of edges $E_{\alpha}=\{e\in E_{\mathbf{G}}\mid f_{\alpha}(e)=1\}$. Our first observation is that $\mathbf{T}=(J,\mu,E_{\alpha})$ is a hyperfinite essential spanning forest.
@@ 4.1 Measures, functions and subsets
Let $G=(J,\mu,E)$ be a graphing and $\mathbf{T}=(J,\mu,F)$ be a spanning subgraphing. The following are equivalent:
  1. (i)
    $\mathbf{T}$ is a hyperfinite spanning forest;
  2. (ii)
    the measure $\alpha_{F}(A)=\eta_{\mathbf{G}}(A\cap F)$ is in $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$.
Similarly, the following are equivalent:
  1. (i’)
    $\mathbf{T}$ is a hyperfinite essential spanning forest;
  2. (ii’)
    the measure $\alpha_{F}(A)=\eta_{\mathbf{G}}(A\cap F)$is in $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$.
@@
[…]
4.2 Exposed points
Let $C$ be a compact, convex subset of a locally convex topological space. We say that $v\in C$ is exposed if there exists $\varphi$ linear, continuous functional such that $v$ is the unique maximizer of $\varphi$ on $C$, that is, $\varphi(v)>\varphi(u)$ for any $u\in C$, $u\neq v$. In this case, we say that $v$ is exposed by the function $\varphi$ and call $\varphi$ a witness function for $v$. It is not difficult to see that exposed points are extreme. In this section we characterize the exposed points of $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ and $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$. We prove that all these points are of the form $\alpha_{F}$ where $F$ is the edge set of some hyperfinite spanning forest, or in other words, all the exposed $f_{\alpha}$ are 00-$1$ valued.
While being extreme is a purely algebraic property, being exposed my depend on the underlying topology. Therefore, first we need to specify which space to embed $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ and $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ into. A natural choice would be using $L^{p}$ with the weak topology for some $p\in[1,\infty)$, or $L^{\infty}$ with the weak topology. As it turns out, the choice of the space does not matter as $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ and $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ are compact and the exposed points are exactly same in all cases. For ease of discussion, we use the space $L^{1}$ with the weak topology.
Theorem 4.2.
Let $f_{\alpha}\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ be exposed by a function $g\in L^{\infty}(E,\eta_{\mathbf{G}})$. Then, $f_{\alpha}$ is 00-$1$ valued $\eta_{\mathbf{G}}$-almost everywhere.
Proof.
We treat $g$ as a weight function on the edges, and our goal is to find a maximum weight hyperfinite essential spanning forest $F$ with respect to this weight function. Recall that $\alpha_{F}$ denotes the measure in $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ defined as $\alpha_{F}(A)=\eta_{\mathbf{G}}(A\cap F)$ for all measurable $A\subseteq E$. We show that there exists an $F$ satisfying $\alpha_{F}(\{g\geq t\})=\rho_{\mathbf{G}}(\{g\geq t\})$ for every $t\in\mathbb{R}$. For any $\beta\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$, in a layer-cake integration this implies
[…]
@@ 4.2 Exposed points
To verify the existence of a required $F$, we run a variant of Kruskal’s greedy algorithm [18] on $\mathbf{G}$, sometimes called Prim’s invasion algorithm. As a tie-breaking rule, we fix an arbitrary Borel ordering of $E$ in advance, and whenever there is a tie among the $g$ values of finitely many edges considered, we choose the maximal one according to the Borel ordering. The algorithm constructs Borel edge sets $F_{1}\subseteq F_{2}\subseteq\dots$, with $F=\bigcup F_{n}$ being the final output. In the first step, every vertex simultaneously chooses the incident edge with the highest $g$-value to be included in $F_{1}$. During the $n^{\mathrm{th}}$ step, every finite component of $F_{n-1}$ adds the edge with the highest $g$-value from the set of edges leaving it if such an edge exists, and thus we obtain $F_{n}$. In case there is no edge leaving a finite component, observe that it is also a component in theoriginal graphing and $F_{n-1}$ is already a spanning tree in it. Also note that infinite components do not add any edges, but it may happen that a finite component adds an edge that connects it to an infinite one.
@@
It is not difficult to see that the sets $F_{n}$ are Borel, cycles cannot be created, and $F$ is essential. We claim that $\eta_{\mathbf{G}}(F)=\rho_{\mathbf{G}}(E)$, so $F$ is hyperfinite by Theorem 2.1. Indeed, set up a payment function as follows. In the beginning, every vertex receives $1$ token. In a general step, every finite connected component gives $1$ token to the edge it chooses. After the addition of the edges, for every finite connected component there is exactly one edge who received $2$ tokens. This edge gives $1$ token back to the component, thus ensuring that every finite component has a token at the beginning of the next step. Should a finite $F$-component become spanning in a finite $\mathbf{G}$-component, and therefore has no use for its token, this token is redistributed equally among the vertices of this component. In the end, vertices in finite $\mathbf{G}$-components paid $1-\frac{1}{|\mathbf{G}_{x}|}$ token, and vertices in infinite $\mathbf{G}$-components paid 1 token. On the other hand, every edge in $F$ received 1 token, hence $\rho_{\mathbf{G}}(E)=\eta_{\mathbf{G}}(F)$ by the Mass Transport Principle. As $F$ is hyperfinite, by Lemma 4.1, we have $\alpha_{F}\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$.
Let $A_{t}=\{g\geq t\}$ and $F_{\geq t}=F\cap A_{t}$. Observe that $F_{t}$ is an essential spanning forest of $A_{t}$. That is, for $\mu$-almost every $x\in J$,
@@
Observations (1) and (2) imply $\alpha_{F}(A_{t})=\rho_{\mathbf{G}}(A_{t})$, completing the proof of the theorem. ∎
The algorithm also proves the following corollary.
Corollary 4.3.
Let $\mathbf{G}$ be a graphing and $F\subseteq E(\mathbf{G})$ be a hyperfinite subforest of $\mathbf{G}$. Then $F$ can be extended to a hyperfinite essential spanning forest of $\mathbf{G}$.
Proof of Corollary 4.3.
Take any Borel ordering of the set $E(\mathbf{G})\setminus F$ and run the same invasion algorithm as in the proof of Theorem 4.2 with the only difference that the starting graph, which was the empty graph in the previous case, is now $F$. ∎
The exposed points of $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ can be characterized in an analogous way.
Theorem 4.4.
Let $f_{\alpha}\in\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ be exposed by a function $g\in L^{\infty}(E,\eta_{\mathbf{G}})$. Then, $f_{\alpha}$ is 00-$1$ valued $\eta_{\mathbf{G}}$-almost everywhere.
Proof.
Run the algorithm from the proof of Theorem 4.2. That gives us a hyperfinite essential spanning forest $F$ with the property that for all $t\in\mathbb{R}$, we have $\alpha_{F}(\{g\geq t\})=\rho_{\mathbf{G}}(\{g\geq t\})$. Now let $F^{\prime}=\{e\in F\mid g(e)\geq 0\}$. We claim that $F^{\prime}$ maximizes the function $g$ on $C$. The proof, again, relies on the cake-layer representation. Take an $f\in\text{\sf mm}_{+}(\rho_{\mathbf{G}})$. We may assume that $f$ is zero on the set $N=\{g<0\}$, otherwise $\varphi_{g}(f)<\varphi_{g}(f{\mathbbm{1}}_{N^{c}})$.
$\varphi_{g}(f)=\int_{N^{c}}g\cdot f\mathop{}\!\mathrm{d}\eta_{\mathbf{G}}=\int_{N^{c}}g\mathop{}\!\mathrm{d}\eta_{f}=\int_{0}^{\infty}\eta_{f}\left(\{g\geq t\}\right)\mathop{}\!\mathrm{d}\lambda(t)\leq$
$\int_{0}^{\infty}\rho_{\mathbf{G}}\left(\{g\geq t\}\right)\mathop{}\!\mathrm{d}\lambda(t)=\int_{0}^{\infty}\alpha_{F^{\prime}}\left(\{g\geq t\}\right)\mathop{}\!\mathrm{d}\lambda(t)=\varphi_{g}({\mathbbm{1}}_{F^{\prime}}).$
[…]
@@
If $f_{\alpha}$ is 00-$1$ valued, then it is exposed – the witness for this is simply choosing $g=f_{\alpha}$. In fact, in this case $f_{\alpha}$ is strongly exposed, i.e. for any sequence $\beta_{n}\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ we have $\varphi_{g}(f_{\beta_{n}})\to\varphi_{g}(f_{\alpha})\Rightarrow||f_{\beta_{n}}-f_{\alpha}||_{p}\to 0$ for all $p\in[1,\infty)$. That is, Theorem 4.2 completely characterizes exposed points of $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$, and a similar statement holds for $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ as well.
@@
By the Klee–Straszewicz theorem [17, Theorem 2.1], exposed points are dense among extreme points of $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ in the weak topology of $L^{\infty}(E,\eta_{\mathbf{G}})$. Moreover, the exposed points determine $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ as $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})=\overline{\textrm{conv}(\exp(\text{\sf bmm}_{+}(\rho_{\mathbf{G}})))}^{w^{*}}$. Again, a similar statement holds for $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ as well.